### v1

# Open quantum battery in the background of a three-dimensional rotating black hole

Abstract
We investigate the charging performance of a quantum battery (QB) coupled to a scalar field under the background of a three-dimensional rotating black hole. It is shown that under Dirichlet boundary conditions, when the QB’s energy level spacing is smaller than the charging amplitude, the black hole rotation enhances the charging performance at finite times, whereas in other parameter regimes, it degrades the charging performance. Notably, as the black hole approaches extremal rotation, charging performance undergoes significant amplification or suppression, depending on the parameter regime. This indicates that the performance of QBs can probe critical properties of black holes. Additionally, regarding the energy flow in QB, it is further demonstrated that the energy extraction from vacuum fluctuations via dissipation, and rotation suppresses the QB’s capacity to extract this energy. Our findings not only advance the relativistic dissipation dynamics of QBs but also propose a novel method to detect black hole rotation and extremal-state transitions.
pacs:
I Introduction
In recent years, the rapid advancement of thermodynamics and quantum information science has intensified research focus on the quantum thermodynamic properties at their intersection Esposito et al. (2009); Campisi et al. (2011); Landi and Paternostro (2021); Pekola and Karimi (2021). An important example is quantum battery Allahverdyan et al. (2004); Alicki and Fannes (2013), a miniaturized device that leverages quantum effects and quantum operations to achieve advantages in energy storage capacity and charging power Yang et al. (2023); Zhang et al. (2024). To investigate how the charging performance of quantum battery can be enhanced by quantum effects, various quantum battery schemes have been studied from both theoretical and experimental perspectives Zhu et al. (2023); Niu et al. (2024); Hu et al. (2022); Quach et al. (2022); Joshi and Mahesh (2022), such as the Sachdev-Ye-Kitaev quantum battery Rossini et al. (2020); Rosa et al. (2020), the cavity quantum battery Crescente et al. (2020); Lu et al. (2021); Zhang et al. (2019); Dou et al. (2022a); Yang et al. (2024); Wang et al. (2024), and the spin-chain quantum battery Kamin et al. (2020); Grazi et al. (2024); Dou et al. (2022b); Ghosh et al. (2020); Rossini et al. (2019). Most early studies focused on closed quantum battery systems, which consist of a quantum battery and charger isolated from environmental interactions Ferraro et al. (2018); Crescente et al. (2020); Zhang et al. (2019). However, real quantum systems always interact with the environment Breuer and Petruccione (2002), causing relaxation and decoherence. For quantum batteries, these effects can significantly impact the charging performance. Therefore, investigating dissipative charging processes in open quantum systems is of critical importance Barra (2019); Bai and An (2020); Zhao et al. (2021); Gherardini et al. (2020); Mitchison et al. (2021); Xu et al. (2021); Tian et al. (2024); Bhattacharyya et al. (2024); Malavazi et al. (2024); Hu et al. (2025); Tian et al. (2025).
On the other hand, as a result of quantum field scattering by spacetime, the geometric properties of spacetime are typically encoded into the field correlation function, which in turn may influence the dynamical evolution of the detector through its interaction with external quantum field Manzano (2020); Benatti et al. (2003); Benatti and Floreanini (2004); Soares et al. (2022); Yu and Zhang (2008); Yu (2011); Kaplanek and Burgess (2020); Kaplanek and Tjoa (2023). Consequently, the study of open quantum systems within the relativistic framework has received extensive attention, including investigations into quantum field and spacetime properties Tian and Jing (2013); Liu et al. (2024c); Du and Mann (2021); Tian et al. (2015); Tian and Jing (2023), harvesting the vacuum entanglement of quantum field Liu et al. (2025b, 2021); Wu et al. (2024), and so on. In addition, this detector-field interaction model has also been extensively employed to investigate the interplay between the thermodynamics of quantum systems and the superposition property of spacetime Henderson et al. (2020); Martín-Martínez and Perche (2023); Foo et al. (2022).
This work investigates the dissipation dynamics of an open quantum battery under the background of a three-dimensional BTZ black hole Banados et al. (1992). The quantum battery is modeled as a two-level system coupled to the external massless scalar field. By focusing on dissipative dynamics, the charging performance of the battery is analyzed to investigate the interplay between the thermodynamics of quantum systems and gravitational effects. Specifically, we investigate the influence of black hole rotation on the charging performance by varying energy level spacing and charging amplitude. Our results show that when the energy level spacing is smaller than the charging amplitude, dissipation under Dirichlet boundary conditions in the extremal black hole scenario significantly enhances both energy storage capacity and charging power. Furthermore, an analysis of energy flow demonstrates that increased black hole rotation speeds reduce the proportion of energy extracted from vacuum fluctuations.
The organization of the paper is as follows. In Sec. II, we introduce the open quantum battery model. In Sec. III, we introduce the rotating BTZ black hole and derive the charging dynamics of an open quantum battery in this background. Based on the derived dynamics expression, we analyze the effect of dissipation on the average energy and charging power of the quantum battery. Finally, our conclusions are given in Sec. IV.
II Dissipation dynamics of open quantum battery
The Hamiltonian of the quantum battery and quantum field reads
$H=H_{\mathrm{B}}+H_{\mathrm{C}}+H_{\mathrm{FI}},$ (1)
where $H_{\mathrm{B}}=\frac{\Delta}{2}\sigma_{z}$ and $H_{\mathrm{C}}=-\frac{A}{2}\sigma_{x}$ denote the free Hamiltonian of the quantum batttery and the charger, respectively. $\Delta$ is the energy level spacing between the ground state $\left|g\right\rangle$ and the excited state $\left|e\right\rangle$ of the quantum battery, $\sigma_{m}(m=x,y,z)$ represents the $m\mathrm{-th}$ Pauli matrix, and $A$ denotes an external classical field coupled with the $\sigma_{x}$ component of the quantum battery Binder et al. (2015); Campaioli et al. (2017); Ferraro et al. (2018). Meanwhile, $H_{\mathrm{FI}}=H_{\mathrm{F}}+H_{\mathrm{I}}$ contains two terms: $H_{\mathrm{F}}$ denotes the free Hamiltonian of the scalar field, and $H_{\mathrm{I}}$ denotes the interaction between the quantum battery and the field. Specifically,
$H_{\mathrm{F}}$ $=$ $\int d^{3}\mathbf{k}\omega_{\mathbf{k}}c_{\mathbf{k}}^{\dagger}c_{\mathbf{k}}\frac{dt}{d\tau},$
$H_{\mathrm{I}}$ $=$ ${\lambda}(\sin\theta\sigma_{z}+\cos\theta\sigma_{x})\otimes\phi[x(\tau)],$ (2)
where $c_{\mathbf{k}}^{\dagger}$($c_{\mathbf{k}}$) is the creation (annihilation) operator of the field mode $\mathbf{k}$ with frequency $\omega_{\mathbf{k}}$, $\tau$ denotes the proper time of the quantum battery, $\lambda$ represents the coupling strength between the quantum battery and the field, and $\theta$ characterizes the weight of the quantum battery-field coupling that along the longitudinal $(z)$ and transverse $(x)$ directions. Note that this interaction can capture both decoherence $(\theta=0)$ and pure dephasing $(\theta=\pi/2)$ process. It is the interaction with the field that will lead to the dissipation of the quantum battery, as shown below.
We assume that the initial state of the quantum battery and field is prepared at $\rho_{\mathrm{tot}}(0)=\rho_{\mathrm{B}}(0)\otimes|0\rangle\langle 0|$, where $\rho_{\mathrm{B}}(0)=\frac{1}{2}(\mathrm{I}+\vec{\mathrm{\textbf{b}}}(0)\cdot\vec{\sigma})$, and $\left|0\right\rangle$ represents the vacuum state of the field. At $\tau>0$, we switch on the charger, then the quantum battery begins to be charged. To learn the dynamics of the quantum battery conveniently, we employ a rotation operation (23) on the spin space of the quantum battery. Correspondingly, the rotated time-dependent reduced density matrix of quantum battery reads
$\tilde{\rho}_{\mathrm{B}}(\tau)=\frac{1}{2}(\mathrm{I}+\vec{\mathbf{a}}(\tau)\cdot\vec{\sigma}).$ (3)
At $\tau=0$, the rotated initial reduced density matrix of the quantum battery is given by $\tilde{\rho}_{\mathrm{B}}(0)=\frac{1}{2}(\mathrm{I}+\vec{{\textbf{a}}}(0)\cdot\vec{\sigma})$, with $a_{1}(0)=b_{1}(0)\cos\Theta+b_{3}(0)\sin\Theta$, $a_{2}(0)=b_{2}(0)$, and $a_{3}(0)=b_{3}(0)\cos\Theta-b_{1}(0)\sin\Theta$, and $\Theta=\mathrm{Arccos}\big{(}{\Delta}/{\sqrt{\Delta^{2}+\mathrm{A}^{2}}}\big{)}.$
Under the assumption of weak-coupling limit between the quantum battery and the field, we apply the Born approximation and the Markov approximation Breuer and Petruccione (2002) to derive the equation of motion of the quantum battery. After that, we obtain the reduced density matrix $\tilde{\rho}_{\mathrm{B}}(\tau)$ yields the Lindblad master equation Manzano (2020); Breuer and Petruccione (2002)
$\frac{\partial\tilde{\rho}_{\mathrm{B}}(\tau)}{\partial\tau}$ $=$ $-i\big{[}\tilde{H}_{\mathrm{eff}},\tilde{\rho}_{\mathrm{B}}(\tau)\big{]}+\sum_{i=\pm,z}\bigg{[}\mathcal{L}_{i}\tilde{\rho}_{\mathrm{B}}(\tau)\mathcal{L}_{i}^{\dagger}-\frac{1}{2}\big{(}\mathcal{L}_{i}^{\dagger}\mathcal{L}_{i}\tilde{\rho}_{\mathrm{B}}(\tau)$ (4)
$+\tilde{\rho}_{\mathrm{B}}(\tau)\mathcal{L}_{i}^{\dagger}\mathcal{L}_{i}\big{)}\bigg{]},$
where $\tilde{H}_{\mathrm{eff}}=\frac{1}{2}\Omega\sigma_{z}=\frac{1}{2}\{\sqrt{\Delta^{2}+A^{2}}+\mathrm{Im}(\Gamma_{+}+\Gamma_{-})\}\sigma_{z}$ is the effective Hamiltonian, and $\Omega$ denotes the effective energy level-spacing of the quantum battery, where the correction term $\mathrm{Im}(\Gamma_{+}+\Gamma_{-})$ is the Lamb shift. The Lamb shift term is far smaller than $\sqrt{\Delta^{2}+A^{2}}$, and usually can be neglected. In addition, it is found that $\mathcal{L}_{+}=\sqrt{\gamma_{+}}\sigma_{+},\mathcal{L}_{-}=\sqrt{\gamma_{-}}\sigma_{-}$ and $\mathcal{L}_{z}=\sqrt{\gamma_{z}}\sigma_{z}$, with $\sigma_{\pm}={\frac{1}{2}}(\sigma_{x}\pm i\sigma_{y})$ and
[…]
where $\Delta\tau=\tau-\tau^{\prime}$, and $W(x(\tau),x(\tau^{\prime}))$ denotes the Wightman function.
Substituting Eq. (3) into Eq. (4), one can solve this mater equation analytically. The corresponding time-dependent state parameters in Eq. (3) are given by
$a_{1}(\tau)$ $=$ $a_{1}(0)\cos(\Omega\tau)e^{-\frac{1}{2}(\gamma_{+}+\gamma_{-}+4\gamma_{z})\tau}-a_{2}(0)\sin(\Omega\tau)$
×$\times e^{-\frac{1}{2}(\gamma_{+}+\gamma_{-}+4\gamma_{z})\tau},$
$a_{2}(\tau)$ $=$ $a_{1}(0)\sin(\Omega\tau)e^{-\frac{1}{2}(\gamma_{+}+\gamma_{-}+4\gamma_{z})\tau}+a_{2}(0)\cos(\Omega\tau)$
×$\times e^{-\frac{1}{2}(\gamma_{+}+\gamma_{-}+4\gamma_{z})\tau},$
$a_{3}(\tau)$ $=$ $a_{3}(0)e^{-(\gamma_{+}+\gamma_{-})\tau}+\frac{\gamma_{+}-\gamma_{-}}{\gamma_{+}+\gamma_{-}}(1-e^{-(\gamma_{+}+\gamma_{-})\tau}).$ (6)
To obtain the time-dependent state $\rho_{\mathrm{B}}(\tau)$, one can rotate the time-dependent state $\tilde{\rho}_{\text{B}}(\tau)$, i.e., $\rho_{\mathrm{B}}(\tau)=\mathcal{R}^{\dagger}\tilde{\rho}_{\text{B}}(\tau)\mathcal{R}$. Note that $\frac{1}{2}(\gamma_{+}+\gamma_{-}+4\gamma_{z})$ denotes the dephasing rate, and $\gamma_{+}+\gamma_{-}$ represents the incoherent relaxation rate, shown in Eq. (II), which are responsible for the dissipation and decoherence. Furthermore, the parameters $\gamma_{\pm},\gamma_{z}$, as shown Eq. (II), depend on the field Wightman function, which is related to the spacetime background and the detector’s world line.
III Quantum battery in rotating BTZ black hole
We investigate a quantum battery coupled to a conformal massless scalar field in the vicinity of a rotating BTZ black hole, with particular attention to how the black hole’s rotation influences the charging performance of the quantum battery through dissipation effects. The line element for the rotating BTZ black hole is expressed as follows Banados et al. (1992)
$ds^{2}=-h(r)dt^{2}+\frac{1}{h(r)}dr^{2}+r^{2}\left[N_{\varphi}(r)dt+d\varphi\right]^{2},$ (7)
where $h(r)=-M+\frac{r^{2}}{\ell^{2}}+\frac{J^{2}}{4r^{2}},$ $N_{\varphi}(r)=-\frac{J}{2r^{2}}$, with $M$ being the mass of the black hole, $J$ being its angular momentum, and $\ell$ being the AdS length. Note that the metric (7) is a vacuum solution of Einstein field equations with a negative cosmological constant. The inner and outer horizons (where $h(r)=0$) of the black hole read
$r_{\pm}^{2}=\frac{1}{2}\left(M\ell^{2}\pm\sqrt{M^{2}\ell^{4}-\ell^{2}J^{2}}\right).$ (8)
Notably, $|J|\leq M\ell$, and in the below analysis, we consider $J\in(0,1)$. When $J/M\ell$ = 0, it indicates a static BTZ black hole. An increase in $J/M\ell$ corresponds to an increase in the rotational velocity of the black hole. The condition $r_{+}=r_{-}$ corresponds to the extremal condition.
To derive the charging dynamics of the quantum battery in the rotating BTZ spacetime, one first needs to known the Wightman function of the scalar field. In this paper, we assume that the scalar field is initially in the Hartle-Hawking vacuum. Since the BTZ spacetime is locally equivalent to the $\mathrm{AdS}_{3}$ spacetime, the Wightman function for the conformal massless scalar field in the rotating BTZ can be represent by Carlip (1995); Lifschytz and Ortiz (1994),
$W_{\mathrm{BTZ}}(x,x^{\prime})$ $=$ $\sum_{n=-\infty}^{\infty}\eta^{n}W_{\mathrm{AdS_{3}}}(x,\Gamma^{n}x^{\prime})=\frac{1}{4\pi\sqrt{2}\ell}\sum_{n=-\infty}^{\infty}\eta^{n}$ (9)
×$\times\bigg{[}\frac{1}{\sqrt{\sigma_{\epsilon}(x,\Gamma^{n}x^{\prime})}}-\frac{\zeta}{\sqrt{\sigma_{\epsilon}(x,\Gamma^{n}x^{\prime})+2}}\bigg{]},$
where $\eta=\pm 1$ denotes the untwisted/twisted scalar field case, and $\Gamma:(t,r,\varphi)\rightarrow(t,r,\varphi+2\pi)$ represents the action of the identification. The parameter $\zeta\in\{-1,0,1\}$ represents the field’s boundary conditions at spatial infinity Avis et al. (1978): Neumann, transparent, and Dirichlet, respectively. Besides, we have
$\sigma_{\epsilon}(x,\Gamma^{n}x^{\prime})^{2}$ $=$ $-1+\frac{\cosh\bigg{[}\frac{r_{+}}{\ell}(\Delta\phi-2\pi n)-\frac{r_{-}}{\ell^{2}}(t-t^{\prime})\bigg{]}}{\big{[}\alpha(r)\alpha(r^{\prime})\big{]}^{-1/2}}$ (10)
$-\frac{\cosh\bigg{[}\frac{r_{+}}{\ell^{2}}(t-t^{\prime})-\frac{r_{-}}{\ell}(\Delta\phi-2\pi n)\bigg{]}}{\big{[}(\alpha(r)-1)(\alpha(r^{\prime})-1)\big{]}^{-1/2}},$
where $\alpha(r)=\frac{r^{2}-r_{-}^{2}}{r_{+}^{2}-r_{-}^{2}}$, and $\Delta\phi=\phi-\phi^{\prime}$.
Here we only consider the untwisted scalar field case. We assume that our quantum battery is co-rotating with the black hole, thus the corresponding world line yields
$t$ $=$ $\frac{\ell r_{+}\tau}{\sqrt{r^{2}-r_{+}^{2}}\sqrt{r_{+}^{2}-r_{-}^{2}}},$
$r$ $=$ $R,$
$\phi$ $=$ $\frac{r_{-}\tau}{\sqrt{r^{2}-r_{+}^{2}}\sqrt{r_{+}^{2}-r_{-}^{2}}}.$ (11)
Substituting the Wightman function (9), together with the quantum battery’s world line (III), into Eq. (II), one can derive
$\gamma_{\pm}$ $=$ $\frac{\lambda^{2}\cos^{2}(\theta-\Theta)}{4}\bigg{[}1-\tanh\left(\pm\frac{\Omega}{2T}\right)\bigg{]}\sum_{n=-\infty}^{\infty}e^{\pm\frac{in\Omega r_{-}}{\ell T}}\bigg{[}P_{\pm\frac{i\Omega}{2\pi T}-\frac{1}{2}}\big{(}\cosh\alpha_{n}^{-}\big{)}-\zeta P_{\pm\frac{i\Omega}{2\pi T}-\frac{1}{2}}\big{(}\cosh\alpha_{n}^{+}\big{)}\bigg{]},$
$\gamma_{z}$ $=$ $\frac{\lambda^{2}\sin^{2}(\theta-\Theta)}{4}\sum_{n=-\infty}^{\infty}\bigg{[}P_{-\frac{1}{2}}\big{(}\cosh\alpha_{n}^{-}\big{)}-\zeta P_{-\frac{1}{2}}\big{(}\cosh\alpha_{n}^{+}\big{)}\bigg{]},$ (12)
where $T$ is the local KMS temperature, which is the temperature experienced by the quantum battery, satisfying $T=T_{H}/\gamma$, with $\gamma=\frac{\sqrt{R^{2}-r_{+}^{2}}\sqrt{r_{+}^{2}-r_{-}^{2}}}{lr_{+}}$ and $T_{H}=\frac{1}{2\pi\ell}\left(\frac{r_{+}^{2}-r_{-}^{2}}{r_{+}}\right)$. Besides, $P_{v}\left(x\right)$ is the associated Legendre function of the first kind, satisfying $P_{-\frac{1}{2}+i\lambda}=P_{-\frac{1}{2}-i\lambda}$, and the definitions
$\cosh(\alpha_{n}^{\pm})={\pm}4\ell^{2}\pi^{2}T^{2}+(1+4\ell^{2}\pi^{2}T^{2})\cosh\left(\frac{2\pi nr_{+}}{\ell}\right),$ (13)
have been used.
With the above specific parameters, $\gamma_{\pm}$ and $\gamma_{z}$, we can analyze how the properties of spacetime influence the dynamics of the quantum battery, thereby affecting its relevant charging performance. Specifically, we will examine the charging performance of the quantum battery in terms of its figures of merit, namely, the average energy stored in the quantum battery during the charging protocol and the average charging power that characterizes the charging speed of the protocol Binder et al. (2015).
III.1 Dynamics of the average energy
In this subsection, we investigate how the rotation property of the rotating BTZ black hole affects the average energy stored in the quantum battery during the charging protocol, through dissipation determined by $\gamma_{\pm}$ and $\gamma_{z}$ shown in Eq. (III). In doing so, we assume the quantum battery is initially prepared in the ground state $|g\rangle$, i.e., $b_{1}(0)=0,b_{2}(0)=0,b_{3}(0)=-1$, and consider the decoherence coupling case ($\theta=0$ shown in Eq. (II)). We can obtain the average energy of the quantum battery and the charger Yang et al. (2023)
$\langle E_{\mathrm{B}}(\tau)\rangle$ $=$ $\frac{\Delta}{2}\bigg{\{}1-\frac{\Delta^{2}}{\Omega^{2}}e^{-\mathcal{K}^{\prime}\tau}-\frac{\Delta}{\Omega}\left(\frac{e^{\Omega/T}-1}{e^{\Omega/T}+1}\right)(1-e^{-\mathcal{K}^{\prime}\tau})$ (14)
$\quad\;\;\ -\frac{A^{2}}{\Omega^{2}}\cos(\Omega\tau)e^{-\mathcal{K}\tau}\bigg{\}},$
$\langle E_{\mathrm{C}}(\tau)\rangle$ $=$ $\frac{A^{2}}{2\Omega}\Bigg{\{}-\frac{\Delta}{\Omega}e^{-\mathcal{K}^{\prime}\tau}-\left(\frac{e^{\Omega/T}-1}{e^{\Omega/T}+1}\right)(1-e^{-\mathcal{K}^{\prime}\tau})$ (15)
$\quad\quad\;\ +\frac{\Delta}{\Omega}\cos(\Omega\tau)e^{-\mathcal{K}\tau}\Bigg{\}},$
where
$\mathcal{K}^{\prime}$ $=$ $\frac{\lambda^{2}\Delta^{2}}{2\Omega^{2}}\sum_{n=-\infty}^{\infty}e^{\pm\frac{i\Omega nr_{-}}{\ell T}}\bigg{[}P_{\frac{i\Omega}{2\pi T}-\frac{1}{2}}\big{(}\cosh\alpha_{n}^{-}\big{)}-\zeta P_{\frac{i\Omega}{2\pi T}-\frac{1}{2}}\big{(}\cosh\alpha_{n}^{+}\big{)}\bigg{]},$
$\mathcal{K}$ $=$ $\frac{1}{2}\mathcal{K}^{\prime}+\frac{\lambda^{2}A^{2}}{2\Omega^{2}}\sum_{n=-\infty}^{\infty}\bigg{[}P_{-\frac{1}{2}}(\cosh\alpha_{n}^{-})-\zeta P_{-\frac{1}{2}}(\cosh\alpha_{n}^{+})\bigg{]}.$ (16)
Note that dissipation effects are reflected in the incoherent relaxation rate $\mathcal{K}^{\prime}$ and the dephasing rate $\mathcal{K}$.
The absence of dissipation between the quantum battery and the field is also a scenario we are concerned about, which corresponds to the coupling constant $\lambda$ = 0 in Eq. (II), when the closed quantum battery is driven by a static protocol and all its all energy is supplied by the charger. In this case, the energy stored in the quantum battery is
$\langle\,E_{\mathrm{B}}(\tau)\rangle=\frac{\Delta}{2}\frac{A^{2}}{\Omega^{2}}\left[1-\cos(\Omega\tau)\right].$ (17)
From Eq. (17), we know that the energy of the quantum battery varies periodically and that the energy of the closed quantum battery reaches its maximum when $\Omega\tau=\pi n\ (n>0\text{ odd numbers})$, i.e., $\langle E_{\mathrm{B}}(\tau)\rangle_{\mathrm{static}}=\Delta\frac{A^{2}}{\Omega^{2}}$. And when $A\gg\Delta$, the stored energy $\langle E_{\mathrm{B}}(\tau)\rangle_{\mathrm{static}}$ approaches its maximum value.
In addition, there is another special scenario in which the quantum battery interacts with the field in the infinite time limit, i.e., $\tau\to\infty$. In this limit, the average energy stored in the quantum battery (14) is expressed as
$\langle\,E_{\mathrm{B}}(\infty)\rangle=\frac{\Delta}{2}\biggl{\{}1-\frac{\Delta}{\Omega}\biggl{(}\frac{e^{\Omega/T}-1}{e^{\Omega/T}+1}\biggr{)}\biggr{\}}.$ (18)
In this limit, the average energy of the quantum battery remains unaffected by exponential decay, as indicated by (III.1). This suggests the boundary conditions have no discernible effect on the energy at this stage. Instead, the average energy is solely determined by the angular momentum $J/M\ell$ and relative weighting between the quantum battery’s energy level spacing $\Delta\ell$ and the charger’s amplitude $A\ell$. The effects of these two factors on the average energy stored in a quantum battery in the asymptotic long-time limit are elucidated Fig. 1.
Figure 1: The average energy $\langle E_{\mathrm{B}}(\infty)\rangle$ stored in the quantum battery, given by Eq. (18), parameterized by the angular momentum $J/M\ell$. The gray surface indicates the maximum average energy $\langle E_{\mathrm{B}}(\tau)\rangle_{\mathrm{static}}$ stored by the closed quantum battery. The parameters are: $\ell=1$, $M=1$, $A\ell=0.1$ and $R=1.01$.
It is demonstrated that the maximum energy stored in a closed quantum battery is solely dependent on the energy level spacing $\Delta\ell$ and charging amplitude $A\ell$. It is noteworthy that no energy can be stored in the closed quantum battery when the ratio of $\Delta\ell$/$A\ell$ = 0. This indicates that although there is a charger providing energy, the energy level spacing $\Delta\ell$ determines the storage capacity, and no energy can be stored in the quantum battery when $\Delta\ell$ = 0. As the $\Delta\ell$/$A\ell$ increase, the maximum energy stored in the closed quantum battery first increases sharply, reaches a peak when the ratio is near one, and then decreases slowly. This means that the charging amplitude $A\ell$ is fixed, the closed quantum battery stores the maximum energy when the energy level spacing $\Delta\ell$ is equal to $A\ell$.
In addition, it is clearly observed that when the energy level spacing $\Delta\ell$ of the quantum battery is larger than the charging amplitude $A\ell$, and as the angular momentum varies, the energy stored in the quantum battery in the long-time limit may exceed the maximum energy of the closed quantum battery. The energy stored in the quantum battery originates from both the charger and dissipation due to rotating BTZ vacuum fluctuations. In other words, the quantum battery can extract energy from the vacuum fluctuations. However, as angular momentum increases beyond approximately 0.9, the stored energy falls below $\langle E_{\mathrm{B}}(\tau)\rangle_{\mathrm{static}}$. This implies that, in the infinite time limit, the black hole’s rotation inhibits the quantum battery’s ability to extract energy from vacuum fluctuations, which in turn weakens its charging performance. When $\Delta\ell$ $<$ $A\ell$, the dissipative quantum battery stores less energy than the closed quantum battery, regardless of angular momentum. This suggests that, in addition to the rotation of the black hole, the weights between $\Delta\ell$ and $A\ell$ affect charging dynamics.
Figure 2: Evolution of the average energy $\langle E_{\mathrm{B}}(\tau)\rangle$ stored in the quantum battery over time. In the first row take $\Delta\ell$ = 0.05, $A\ell$ = 0.1. In the second row take $\Delta\ell$ = 0.15 and $A\ell$= 0.1. Here we set the mass $M=1$, the radial coordinates of quantum battery $R=1.01$ and coupling strength $\lambda\sqrt{\ell}=0.2$. The black dotted line indicates the maximum average energy of the quantum battery when the coupling strength $\lambda\sqrt{\ell}=0$.
In Fig. 2 the average energy $\langle E_{\mathrm{B}}(\tau)\rangle$ of the quantum battery shown in (14) is plotted as a function of the charging time. We fix the charging amplitude as $A\ell=0.1$, the rotating BTZ black hole mass as $M=1$, and the location of the quantum battery as $R=1.01$. Representative examples of different energy-level spacing of the quantum battery $\Delta\ell=0.05$ (Fig. 2 (a), (b)), and $\Delta\ell=0.15$ (Fig. 2 (c), (d)) have been chosen. We also plot the maximum average energy stored in the closed quantum battery driven by a static driving protocol (corresponding to $\lambda\sqrt{\ell}=0$ case in the interacting Hamiltonian given in Eq. (II)) as a reference limit (see black dotted curves), which is given by $\langle E_{\mathrm{B}}(\tau)\rangle_{\text{static}}=\Delta\frac{A^{2}}{\Omega^{2}}$.
When the energy level spacing $\Delta\ell$ of the quantum battery is smaller than the charging amplitude $A\ell$, under Dirichlet boundary conditions with dissipation (Fig. 2(a)), stored in the quantum battery exhibits damped oscillations over time, with amplitude decaying exponentially due to incoherent relaxation and dephasing rates Eq. (III.1). Dissipation decreases the maximum average energy in an open quantum battery compared to a closed one. Higher angular momentum leads to greater maximum average energy. In limited charging time, increasing the black hole’s rotational speed via dissipation can boost the average energy stored in the quantum battery. However, in the infinite time limit, the average energy converges to a constant, which is given by Eq. (18). In addition, higher angular momentum leads to lower asymptotic average energy in the quantum battery. For transparent boundaries, it reduces both the maximum and asymptotic average energy during charging, unlike Dirichlet boundaries. Thus, increasing black hole rotation decreases stored energy under transparent boundary conditions.
If $\Delta\ell>A\ell$, we can see from Fig. 2(c) and Fig. 2(d) that the average energy exhibits a quite different behavior compared with the $\Delta\ell<A\ell$ case shown in the Fig. 2(a) and Fig. 2(b). Notably, the maximum and asymptotic average energy stored in the open quantum battery can exceed that of the closed quantum battery during the charging process. This indicates the quantum battery extracts energy not only from the charger but also from external vacuum fluctuations in rotating BTZ spacetime, similar to the non-rotating BTZ black hole Tian et al. (2025). Additionally, lower angular momentum increases the stored average energy, meaning black hole rotation suppresses energy extraction and weakens charging performance. This applies to both Dirichlet and transparent boundary conditions. Note: whether the open quantum battery outperforms the closed one depends on the black hole’s rotation.
The behavior of the average energy also depends on the boundary condition of the quantum field at the spatial infinity very much. Compared Fig. 2(a) with Fig. 2(b), we can find that the boundary condition may induce absolutely opposite behavior of the average energy when varying the angular momentum for the $\Delta\ell<A\ell$ scenario. In these two boundary condition cases, they have different maximum and asymptotic evolutionary dynamics. However, for the $\Delta\ell>A\ell$ case, the transparent boundary condition and the Dirichlet boundary condition lead to the same behavior of the average energy via the charging time by varying the angular momentum, while quantitatively different. Moreover, the greater the angular momentum $J/M\ell$, the more pronounced the effect of the boundary conditions on the dynamical behavior of the average energy. However, although the average energy stored in the quantum battery during finite-time charging exhibits different evolution for different boundary conditions, in the infinite-time limit (18), the stored energy depends only on the rotation of the BTZ black hole and is independent of the boundary conditions.
Figure 3: Variation of the maximum average energy $\langle\,E_{\mathrm{B}}\rangle_{\text{max}}$ of the quantum battery as a function of angular momentum $J/M\ell$ when $\Delta\ell=0.05$ and $A\ell=0.1$. The other parameters are: $\ell=1$, $M=1$ and $R=1.01$.
The maximum energy stored in a quantum battery is also a key metric for evaluating its performance
$\langle\,E_{\mathrm{B}}\rangle_{\text{max}}=\max_{\tau}[\langle E_{\mathrm{B}}(\tau)\rangle]=\langle E_{\mathrm{B}}(\tau_{E})\rangle,$ (19)
where $\tau_{E}$ is the time when the average energy reaches its maximum value.
In Fig. 3 we plot the maximum average energy $\langle\,E_{\mathrm{B}}\rangle_{\text{max}}$ stored in the quantum battery as a function of the angular momentum. It is found that with the increase of angular momentum, the maximum average energy increases monotonously for the Dirichlet boundary condition case. However, for the transparent boundary case the maximum average energy decreases monotonously when the angular momentum increases. Furthermore, for the extreme case of the rotating BTZ black hole, i.e., the angular momentum approaches to $1$, the average energy, for both the Dirichlet and the transparent boundary cases, is quite sensitive to the variation of the angular momentum. This means that as the black hole approaches the extremal limit, even small changes in the rotational speed can trigger abrupt changes in the amount of maximum average energy stored in the quantum battery.
III.2 Effect of dissipation on charging power
In this subsection, we employ the average charging power to characterize the “charging speed” Binder et al. (2015), which is defined as the ratio between the average deposited energy $\langle E_{\mathrm{B}}(\tau)\rangle$ and the time $\tau$ required to complete the procedure,
$P_{\mathrm{B}}(\tau)=\frac{\langle E_{\mathrm{B}}(\tau)\rangle}{\tau}.$ (20)
Correspondingly, the maximum value of $P_{\mathrm{B}}(\tau)$ is called the maximal power, given by
$P_{\text{max}}\equiv\max_{\tau}[P_{\mathrm{B}}(\tau)]\equiv P_{\mathrm{B}}(\tau_{P}),$ (21)
where $\tau_{P}$ is the time when the average charging power reaches its maximum value. We are interested in how the rotation property of the rotating BTZ black hole affects the charing performance of the quantum battery from the perspective of the average charging power.
Figure 4: The average charging power $P_{\mathrm{B}}(\tau)$ versus the charging time $\tau/\ell$ for different values of the angular momentum $J/M\ell$. (a)-(b): $\Delta\ell$ = 0.05, $A\ell$ = 0.1; (b)-(c) $\Delta\ell$ = 0.15, $A\ell$ = 0.1.The other parameters are: $\ell=1$, $M=1$, $R=1.01$ and $\lambda\sqrt{\ell}=0.2$.
In Fig. 4 we plot the average charing power as a function of the charging time with varying the angular momentum while fixing other parameters. When $\Delta\ell<A\ell$, the average charing power increases with angular momentum under the Dirichlet boundary conditions but decreases under the transparent boundary conditions. This indicates that black hole rotation enhances charging performance for Dirichlet boundaries while weakening it for transparent boundaries. These are shown in Fig. 4 (a) and (b). When $\Delta\ell>A\ell$, we can see from Fig. 4 (c) and (d) that for both the Dirichlet and transparent boundary conditions, the average charing power is weakened by the angular momentum, which is quite different from the $\Delta\ell<A\ell$ case. This means that, in the case $\Delta\ell>A\ell$, the boundary conditions do not affect the degradation of the charging performance of the quantum battery caused by the rotation of the black hole.
Figure 5: Variation of the maximum charging power $P_{\text{max}}$ with the angular momentum $J/M\ell$. The first row shows $\Delta\ell$ = 0.05, $A\ell$ = 0.1; the second row shows $\Delta\ell$ = 0.15, $A\ell$ = 0.1. The other parameters are: $\ell=1$, $M=1$, $R=1.01$ and $\lambda\sqrt{\ell}=0.2$.
In Fig. 5, we show how the maximum charging power $P_{\max}$ is affected by the angular momentum. When $\Delta\ell<A\ell$, it is shown in Fig. 5 (a) that the maximum charging power increases monotonously with the increase of the angular momentum. Near the extreme angular momentum, the maximum charging power becomes highly sensitive to small changes in the black hole’s rotational speed, significantly improving the quantum battery’s charging performance under the Dirichlet boundary condition. However, under the transparent boundary condition, the maximum charging power decreases monotonically with increasing angular momentum and becomes more sensitive near the extreme condition, as shown in Fig. 5 (b). When $\Delta\ell>A\ell$, the maximum charging power decreases monotonously with the increase of the angular momentum for both the Dirichlet and transparent boundary condition cases, as shown in Fig. 5 (c) and (d). However, the maximum charging power for the Dirichlet boundary condition case decrease faster than that for the transparent boundary condition case when the angular momentum is relatively small, while decreases more slowly when the angular momentum is near the extreme case. This further reveals the distinct effects of boundary conditions in the context of a quantum battery whose charging performance is influenced by the rotation of a black hole.
III.3 Energy flow in quantum battery
From the above analysis, we find that whether the open quantum battery outperforms the closed one depends on the ratio of energy level spacing to charging amplitude. To better understand this charging performance physics, we investigate the flow of average energy during the charging process, as shown in Fig. 6. It is found that when $\Delta\ell<A\ell$, curves associated with the quantum battery and the charger are not specular with respect to the $x$-axis, this reflects that a certain amount of energy is dissipated into the reservoir. Similarly, the curves associated with the quantum battery and the charger are also not specular with respect to the $x$-axis when $\Delta\ell>A\ell$. And the average energy $\langle E_{\mathrm{B}}(\tau)\rangle$ change associated with the quantum battery is greater than the average energy $\langle E_{\mathrm{C}}(\tau)\rangle$ change associated with the charger, which reflects the fact that the average energy of the quantum battery is not only obtained from the charger but also extracted from the vacuum fluctuations of the quantum fields in rotating BTZ black hole.
Moreover, from the dynamics of the average energy evolution, it is evident that the dissipation plays a dominant role in the overall charging process at this stage, as shown in Fig. 6(c) and (d). It is worth noting that according to the variation of the average energy dynamics curves, the quantum battery would extract less energy from the vacuum fluctuations when the angular momentum is larger compared to when the angular momentum is smaller. This means that the rotation of the BTZ black hole inhibits the quantum battery from extracting energy from the vacuum fluctuations.
Figure 6: Evolution of average energy $\langle E\rangle$ over time. We take $\Delta\ell$ = 0.05, $A\ell$ = 0.1 in (a)-(b) and $\Delta\ell$ = 0.15, $A\ell$ = 0.1 in (c)-(d). Solid lines are associated with the quantum battery, and dashed lines are associated with the charger. Other parameters are the same as those in the graphs above.
IV Conclusions
In summary, we studied the dissipative dynamics of a quantum battery coupled to the massless scalar field of a rotating BTZ black hole. The effect of black hole rotation on the charging performance is explored, focusing on decoherence coupling regime. We find that the boundary conditions and the ratio between energy level spacing and charging amplitude determine how black hole rotation impacts the charging performance. When energy level spacing is smaller/larger than the charging amplitude, black hole rotation can enhance/weaken performance. Increasing rotational velocity strengthens the influence of boundary conditions on average energy dynamics. Moreover, energy extraction from vacuum fluctuations via dissipation in curved spacetime is demonstrated, with black hole rotation reducing extractable energy. In specific cases, such as extremal black holes, rotation can significantly enhance/weaken the charging performance. This reveals that quantum batteries in strong gravitational fields can probe critical black hole properties. Our study elucidates a new mechanism where black hole spin regulates quantum energy storage through vacuum fluctuations.
Acknowledgements.
This work was supported by the National Natural Science Foundation of China under Grants No.12475051, No.12374408, and No.12035005; the science and technology innovation Program of Hunan Province under grant No.2024RC1050; the Natural Science Foundation of Hunan Province under grant No.2023JJ30384; and the innovative research group of Hunan Province under Grant No.2024JJ1006. ZT was supported by the scientific research start-up funds of Hangzhou Normal University: 4245C50224204016.
Appendix A Charging performance indicators
To describe the charging dynamics of quantum battery, the time evolution of spin component $\langle\sigma_{m}(\tau)\rangle$ is defined as
$\langle\sigma_{m}(\tau)\rangle=\mathrm{Tr}[\rho(\tau)\sigma_{m}],$ (22)
where $\rho(\tau)=\mathrm{Tr}_{\mathrm{F}}[\rho_{\mathrm{tot}}(\tau)].$
We investigate the dynamical evolution of different subparts using the unitary rotation of the spin space
$\mathcal{R}=e^{-i\frac{\Theta}{2}\sigma_{y}},$ (23)
where the phase factor $\Theta$ is selected as the value to project $H_{\mathrm{B}}$ + $H_{\mathrm{C}}$ only along the $z$-axis. Following a unitary rotation, one obtains
$\tilde{H}_{\mathrm{B}}=\mathcal{R}H_{\mathrm{B}}\mathcal{R}^{\dagger}=\frac{\Delta}{2}(\cos\Theta\sigma_{z}+\sin\Theta\sigma_{x}),$ (24)
$\tilde{H}_{\mathrm{C}}=\mathcal{R}H_{\mathrm{C}}\mathcal{R}^{\dagger}=\frac{A}{2}(\sin\Theta\sigma_{z}-\cos\Theta\sigma_{x}),$
$\tilde{H}_{\mathrm{I}}=\mathcal{R}H_{\mathrm{I}}\mathcal{R}^{\dagger}=\lambda(\sin(\theta-\Theta)\sigma_{z}+\cos(\theta-\Theta)\sigma_{x})\otimes\phi(x(\tau)).$
Moreover, by combining $\tilde{H}_{\mathrm{B}}$ and $\tilde{H}_{\mathrm{C}}$ in Eq. (24), we can calculate both the average energy $\langle H_{\mathrm{B}}(\tau)\rangle$ of the quantum battery and the average energy $\langle H_{\mathrm{C}}(\tau)\rangle$ of the charger expressed in Eq. (II)
$\langle H_{\mathrm{B}}(\tau)\rangle=\frac{\Delta}{2}\left[a_{3}(\tau)\cos\Theta+a_{1}(\tau)\sin\Theta\right],$ (25)
$\langle H_{\mathrm{C}}(\tau)\rangle=\frac{A}{2}\left[a_{3}(\tau)\sin\Theta-a_{1}(\tau)\cos\Theta\right].$
The foregoing equations clearly show that the charge dynamics depend on the spacetime nature (II), and in turn the charging dynamics can be used to learn about the spacetime nature.
The charging dynamics of the quantum battery can be then determined by studying the energy changes between the different subparts at time $\tau$
$\langle E_{j}(\tau)\rangle=\langle H_{j}(\tau)\rangle-\langle H_{j}(0)\rangle,$ (26)
where $j$ = B, C, FI. For $\tau>0$, the driving is assumed to be static, implying $\dot{H}=0$ and the system’s energy balance $\langle E_{\mathrm{B}}(\tau)\rangle+\langle E_{\mathrm{C}}(\tau)\rangle+\langle E_{\mathrm{FI}}(\tau)\rangle=0.$ The formula applies to arbitrary drive amplitude and coupling strength.